OpenStudy (anonymous):

Solve -2x + 1 less than or greater to -7 and describe the graph of the solution. x less than or greater to 4; closed circle on 4, shading to the left x greater than or equal to 4; closed circle on 4, shading to the right x greater than or equal to 4; open circle on 4, shading to the right x less than or greater to 4; open circle on 4, shading to the left

OpenStudy (texaschic101):

-2x + 1 <= -7 -2x <= -7 - 1 -2x <= -8 (when dividing by a negative, it changes the inequality sign x >= 4 closed circle on 4,shading to the right

OpenStudy (texaschic101):

any questions ? Quick lesson : Solve inequalities like equations by getting a single variable on one side. One rule to remember with inequalities is that when dividing/multiplying by a negative, changes the inequality sign. If there is an equal sign involved, there is a closed circle. If there is no equal sign, then it is an open circle. If it is greater then, shading goes to the right. If it is less then, shading goes to the left.

OpenStudy (anonymous):

Can you help me with this please!!!!! Choose the inequality that matches the graph below. number line with an open circle on -2 and a closed circle on 4 with shading in between. x > -2 or x less than or greater to 4 x > -2 and x less than or greater to 4 x greater than or equal to -2 and x < 4 x < -2 or x greater than or equal to 4

OpenStudy (texaschic101):

open circle on -2 means that it does not contain an equal sign. The closed circle on 4 means there is an equal sign. Shading in between means that x is greater then -2 and less then 4. However,there is an equal sign connected with the 4, so open circle on -2, closed circle on 4, shading in between your answer is B.

OpenStudy (anonymous):

Thanks:)

OpenStudy (anonymous):

Do you think you can help me with this too? Choose the inequality that matches the graph below. number line with an open circle on -3 with shading to the left and a closed circle on 5 with shading to the right. x less than or greater to -3 and x > 5 x < -3 or x greater than or equal to 5 x < -3 or x > 5 x < -3 and x greater than or equal to 5